OpenStudy (anonymous):

What is the equation of the line that passes through (–2, 3) and is parallel to 2x + 3y = 6? A. B. C. D.

OpenStudy (anonymous):

Right answer 2x+3y = 5; So normal vector have parellel thus equation of the new line is: 2x +3y = q q determine by point (–2, 3) as q=2*(–2)+3*3=5